OpenStudy (anonymous):

Write the vertex form of the parabola with the focus (-68/8,-7) and Directrix x=-65/8

OpenStudy (anonymous):

A parabola is the locus of all points equidistant from a point called focus and a line called directrix.

OpenStudy (anonymous):

Applying the above gives \[\left(x+\frac{68}{8}\right)^2+(y +7)^2=\left(x+\frac{65}{8}\right)^2 \]

OpenStudy (anonymous):

\[ (y+7)^2=\left(x+\frac{65}{8}\right)^2-\left(x+\frac{68}{8}\right)^2 \\ (y+70)^2=-\frac{1}{4} \left(x+\frac{133}{16}\right ) \]

OpenStudy (anonymous):

See the graph.